Fe2O3+3CaS = Fe2S3+3CaO

How many grams of Fe2O3 are needed to produce 51.9 grams of CaO?


1
Expert's answer
2021-03-04T06:42:48-0500

Solution:

Fe2O3 + 3CaS = Fe2S3 + 3CaO

x g 51,9 g

(Fe2O3) = 56*2+16*3= 160g/mol

(CaO)=40+16=56g/mol

Xmol(CaO)= 51,9/56= 0,96

Xmol(Fe2O3)=0,92/3=0,308

Xg(Fe2O3)=160*0,308/1=49,28

Answer:

m(Fe2O3)=49,28g